# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v1

@@ IV Conclusion
As noted above, we seek to understand the number of masked defectives (rather than all masked items). Let $\widetilde{M}_{1}^{\prime}$ and $\widetilde{M}^{\prime}$ respectively denote the number of masked defective items and total masked items respectively between the first and second steps [^7] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. of the generation of $S$ described above. By the same reasoning as above, we have $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$. Moreover, we have $k-2\sqrt{k}\log\tilde{n}\leq\lvert S^{\prime}\rvert\leq k$ with probability $1-o(1)$ due to the multiplicative Chernoff bound (Appendix -A), which implies that $\widetilde{M}_{1}$ equals the number of masked defectives under the combinatorial prior up to $o(1)$ probability events. Combining these observations, it is sufficient to show a high-probability lower bound on $\widetilde{M}_{1}^{\prime}$.
Recall that $(\widetilde{M}_{1}^{\prime}\mid\widetilde{M}^{\prime}=\tilde{m}^{\prime})\sim\mathrm{Bin}(\tilde{m}^{\prime},\tilde{q}^{\prime})$, since under the i.i.d. prior the posterior distribution of a masked item matches its prior [29, Lemma 3.1]. Under the high probability condition on $\widetilde{M}^{\prime}$ stated in (97), the conditional mean of $\widetilde{M}_{1}^{\prime}$ is lower bounded as follows:
@@
with probability $1-o(1)$ due to binomial concentration bounds (Appendix -A). Since $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$, this then implies that
@@
Writing $\tilde{n}=k^{\frac{1}{\theta^{\prime}}}$ and applying some algebraic simplifications, (102) yields that with probability $1-o(1)$, the number of masked defectives $\widetilde{M}_{1}$ is at least
@@
An outline of achievable approximate recovery rates using DD/COMP is given in [8, Section 5.1] for the case that $\eta^{-},\eta^{+}$ are fixed constants $(0,1)$. However, we are unaware of any works giving full proofs, and moreover, the result that we seek in Theorem 1 also permits smaller values of $\eta^{-},\eta^{+}$, which is non-trivial to achieve. Thus, we provide a detailed proof for completeness, despite amounting to relatively straightforward extensions of existing work.
@@
Recalling that $L=\frac{T\log 2}{k}$, we can apply this result with ${\mathcal{M}}=S$ and $\alpha=\log 2$, so that $W^{({\mathcal{M}})}$ corresponds to the number of positive tests. Hence, $W^{({\mathcal{M}})}=T_{1}$ concentrates around $\frac{T}{2}$ (i.e., $|T_{1}-\frac{T}{2}|\leq\delta T$ with probability $1-o(1)$). With this established, we can now proceed to bound the probability that a non-defective item only appears in positive tests. Fix some $i\in[n]\setminus S$; given $T_{1}$, each test that $i$ is drawn to be included in is positive with probability $\frac{T_{1}}{T}$. Thus we can bound the probability of $i$ only being included in positive tests as follows:
@@
where $L_{i}$ is the number of tests in which $i$ appears and no other PD items appear. If we can show that this is the case, then this implies that the number of these false positive items is at most $\eta^{-}k$ with probability $1-o(1)$ due to Markov’s inequality (since the above sum is the expected number of such items).
@@
where (133) holds by the assumption that $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for some $\lambda\in(0,1)$, provided that $\delta$ is sufficiently small. Denoting the number of items $i\in S$ such that $L_{i}=0$ by $Z$, it holds that $\mathbb{E}Z=\Lambda$. Using the aforementioned bound on $\Lambda$ and Markov’s inequality, we thus have the following high probability bound on there being $\eta^{-}k$ or more such items:
$P_{\mathrm{e}}^{-}$ $=\mathbb{P}(Z\geq\eta^{-}k)=o(1).$ (134)

### v3

@@ IV Conclusion
As noted above, we seek to understand the number of masked defectives (rather than all masked items). Let $\widetilde{M}_{1}^{\prime}$ and $\widetilde{M}^{\prime}$ respectively denote the number of masked defective items and total masked items respectively between the first and second steps [^7] Specifically, we use this terminology to mean the number of defectives that would be masked if $S^{\prime}$ were the defective set. of the generation of $S$ described above. By the same reasoning as above, we have $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$. Moreover, we have $k-2\sqrt{k}\log\tilde{n}\leq\lvert S^{\prime}\rvert\leq k$ with probability 1o(1)1-o(1) due to the multiplicative Chernoff bound (Appendix -A), which implies that $\widetilde{M}_{1}$ equals the number of masked defectives under the combinatorial prior up to $o(1)$ probability events. Combining these observations, it is sufficient to show a high-probability lower bound on $\widetilde{M}_{1}^{\prime}$.
[…]
@@
with probability 1o(1)1-o(1) due to binomial concentration bounds (Appendix -A). Since $\widetilde{M}_{1}^{\prime}\leq\widetilde{M}_{1}$, this then implies that
@@
Writing $\tilde{n}=k^{\frac{1}{\theta^{\prime}}}$ and applying some algebraic simplifications, (102) yields that with probability 1o(1)1-o(1), the number of masked defectives $\widetilde{M}_{1}$ is at least
@@
An outline of achievable approximate recovery rates using DD/COMP is given in [8, Section 5.1] for the case that $\eta^{-},\eta^{+}$ are fixed constants in $(0,1)$. However, we are unaware of any works giving full proofs, and moreover, the result that we seek in Theorem 1 also permits smaller values of $\eta^{-},\eta^{+}$, which is non-trivial to achieve. Thus, we provide a detailed proof for completeness, despite amounting to relatively straightforward extensions of existing work.
@@
Recalling that $L=\frac{T\log 2}{k}$, we can apply this result with ${\mathcal{M}}=S$ and $\alpha=\log 2$, so that $W^{({\mathcal{M}})}$ corresponds to the number of positive tests. Hence, $W^{({\mathcal{M}})}=T_{1}$ concentrates around $\frac{T}{2}$ (i.e., $|T_{1}-\frac{T}{2}|\leq\delta T$ with probability 1o(1)1-o(1)). With this established, we can now proceed to bound the probability that a non-defective item only appears in positive tests. Fix some $i\in[n]\setminus S$; given $T_{1}$, each test that $i$ is drawn to be included in is positive with probability $\frac{T_{1}}{T}$. Thus we can bound the probability of $i$ only being included in positive tests as follows:
@@
where $L_{i}$ is the number of tests in which $i$ appears and no other PD items appear. If we can show that this is the case, then this implies that the number of these false positive items is at most $\eta^{-}k$ with probability 1o(1)1-o(1) due to Markov’s inequality (since the above sum is the expected number of such items).
@@
where (133) holds by the assumption that $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for some $\lambda\in[0,1)$, provided that $\delta$ is sufficiently small. Denoting the number of items $i\in S$ such that $L_{i}=0$ by $D$, it holds that $\mathbb{E}D=\Lambda$. Using the aforementioned bound on $\Lambda$ and Markov’s inequality, we thus have the following high probability bound on there being $\eta^{-}k$ or more such items:
$P_{\mathrm{e}}^{-}$ $=\mathbb{P}(D\geq\eta^{-}k)=o(1).$ (134)